A solution of 53% pesticide is to be mixed with a solution of 21% pesticide to form 128 liters of a 40% solution. How many liters of the 53% solution must be used?
Answer by jorel1380(3719)

128x.4=51.2
p+q=128
.53p+.21q=51.2
53p+21q=5120
21p+21q=2688
32p=2432
p=76
q=52
You need 76 liters of 53% solution and 52 liters of 21% solution..